Please help me….my daughter needs to solve the equation below and I have gotten very rusty with my Algebra…Thank you so much
“/” means divide
3(x+5)/6= x+1
Thank you again

Answer by jojo14344(1513)

Absolutely!
3(x+5)/6=x+1
(3x+15)/6=x+1
3x+15=6x+6
6x-3x=15-6
3x=9
x=3, cool!